During the past 8 years, Beef Wellington Cattle Company's common stock dividends have grown from $2.00 to $3.19. Estimate the compound annual dividend growth rate over the 8-year period.
Answer #1
CAGR = (Ending Balance / Beginning Balance)^(1/n)   - 1
CAGR = (3.19/2)^(1/8) - 1   = 1.06 - 1 = 0.06 OR 6%